OpenStudy (anonymous):

What is the distance between points (12, 4) and (18, –1)? Round to the nearest tenth. 7.8 6.1 4.7 9.9

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

hartnn (hartnn):

d= \(\sqrt{6^2+5^2}=7.81\)